Without solving each equation, determine the nature of the roots.
a. 3y^2-10y+1=0
b. 2w^2+6w+10=0

Answer by stanbon(48558)

Without solving each equation, determine the nature of the roots.
a. 3y^2-10y+1=0
b^2-4ac = 100-4*3*1>0 so there are two unequal Real Number roots.
---------------------
b. 2w^2+6w+10=0
b^2-4ac = 36-4*2*10<0 so there are two unequal complex number roots.
=================
Cheers,
Stan H.